# Understanding Trends, Patterns, and Dynamics in Global Acquisitions: A Network Perspective

### v2

@@ Abstract
Studying acquisitions offers invaluable insights into startup trends, aiding informed investment decisions for businesses. However, the scarcity of studies in this domain prompts our focus on shedding light in this area. Employing Crunchbase data, our study delves into the global network of company acquisitions using diverse network analysis techniques. Our findings unveil an acquisition network characterized by a primarily sparse structure comprising localized dense connections. We reveal a prevalent tendency among organizations to acquire companies within their own country and industry. Furthermore, our temporal analysis indicates a growth in network communities over time, accompanied by a trend toward a sparser network. Through centrality metrics computation in the cross-city acquisition network, we identify New York, London, and San Francisco as pivotal and central hubs in the global economic landscape. Finally, we show that the United States, United Kingdom, and Germany are predominant countries in international acquisitions.
@@ 1 Introduction
Crunchbase [^1] https://www.crunchbase.com, PitchBook [^2] https://pitchbook.com, and VentureSource [^3] https://www.venture-source.com are among the most popular platforms providing extensive datasets, collecting information about global firms and their transactions, making them valuable data sources for evaluating acquisitions [7]. However, Crunchbase stands out as the most complete and comprehensive database among them, covering the greatest number of entries and features [8]. Additionally, this platform offers detailed information regarding acquisitions, including type, date, and price, making it a useful data source for analyzing various facets of these transactions. Therefore, we leverage this insightful data for our research to provide a holistic view of company acquisitions.
In this study, we construct and investigate the acquisition network of companies, exploring its projections onto acquiring and acquired firms, cities, and countries using Crunchbase data. We employ a network science approach in our analyses and interpret various structural features of these networks to gain a deeper understanding of dynamics and trends in acquisitions over time.
@@
  • What features influence the interconnections between organizations within the acquisition network?
  • How has the structure of the acquisition network evolved over time?
  • Which cities or countries serve as global economic hubs in terms of acquisitions?
@@
    We calculate and interpret structural features such as transitivity, average clustering coefficient, and the number of communities for these networks.
@@
The rest of the paper is structured as follows. In the “Related work” section, we explore previous studies based on Crunchbase data and juxtapose their scope and perspectives with ours. In the ”Methods and materials” section, we provide an explanation of our data and its features along with the definition of metrics that we calculate in this study. In the ”Results and discussion” section, we describe our analyses and findings and then interpret them. Finally, we conclude the paper and provide directions for future work in the ”Conclusion” section.
@@ 2 Related work
Different economic and managerial studies have utilized Crunchbase data since its foundation in 2007 to gain insights into company interactions, market trends, and strategic behaviors [9, 10]. They have employed various perspectives, focused on different sectors, geographic areas and time brackets, and utilized a wide range of data analysis techniques to address their research inquiries. We provide a general classification of these research works below to offer a more comprehensive view of previous scholarly work in this field.
2.1 Business success prediction
Several studies have utilized Crunchbase data to forecast firms’ success using various methodologies. In [11], the authors constructed machine learning models leveraging demographic, geographical, and fundamental company information obtained from Crunchbase as predictors to gauge the success of young companies, defined by their completion of a series B round, undergoing an IPO, or being acquired. Their findings revealed that the gradient boosting algorithm surpassed other classifiers in accuracy for this prediction task, demonstrating its potential application in real-world scenarios. Te et al. [12] enhanced Crunchbase data with information sourced from LinkedIn and utilized gradient boosting models to predict the successful attainment of Series A funding as a pivotal marker of startup success. Their research demonstrated that integrating this additional data substantially improved model performance, resulting in a higher AUC. In the study conducted by Kim et al. [13], they investigated the influence of industry-specific features on predicting startup success. Employing various machine learning models, they highlighted the significance of these characteristics. Another group of researchers employed a network-based approach to predict the future economic success of young firms, emphasizing the pivotal role of startups’ positions within their ecosystem [14].
2.2 Investment network analysis
Some studies employed network analysis techniques to examine investor behavior utilizing Crunchbase data. Lian and Yuan constructed a social network comprising individuals and companies using this dataset, calculating its characteristics to forecast investors’ funding behavior. Their findings revealed that proximity between investors and companies enhances the likelihood of investment, whereas a higher count of shared connections exhibits a negative impact [15]. The authors of [16] used network alignment methods and proposed intersect proportions to distinguish between investor commitments to startups and entrepreneurs. In [17], the authors introduce an investment behavior prediction model by constructing heterogeneous networks involving various entities within Crunchbase data. They demonstrated the effectiveness of their model as a similarity measure between investors and companies. Holicka and Vinodrai [18] concentrated on blockchain firms and examined city-region level networks to evaluate investment interactions. Their study unveiled that investments in tech startups primarily originate from a select few global cities.
2.3 Industry-specific analysis
Some studies have concentrated on analyzing market trends within specific industries using Crunchbase data. For instance, the authors of [19] explored the evolution of software startup ecosystems. They scrutinized various stages along this trajectory, elucidating the requisites and motives for each developmental phase. Halminen et al. [20] investigated the factors influencing the allocation of venture capital funding for healthcare AI companies using Crunchbase data. Their findings indicated that the presence of research and pharmaceutical organizations as clients positively correlates with increased funding for these companies. In [21], researchers explored the origins of manufacturing startups, examining features by integrating Crunchbase data on experts and managers at these firms with information from their LinkedIn-published CVs. Their findings highlighted the significance of prior experience and academic or tech backgrounds among the teams within these young firms. Eickhoff et al. [22] utilized Crunchbase data to classify FinTech business models. Another group of scholars harnessed this database to assess investor funding preferences specifically within cloud computing startups [23]. They emphasized that investors in this domain show a preference for funding startups offering services in crucial, emerging, and rapidly expanding IT sectors.
2.4 Merger and acquisition analysis
@@
Only a few studies have sought to examine merger and acquisition patterns using Crunchbase data. In [24], the authors constructed temporal networks representing international mergers and acquisitions from this dataset, analyzing both the topological and geographical characteristics of these networks. Additionally, they conducted a comparative analysis between these networks and the international trade network, aiming to discern differences and similarities, thus gaining valuable insights. Guo et al. [25] concentrated on Chinese companies while investigating merger and acquisition networks. They computed various centrality measures to scrutinize M&A behaviors and constructed multiple linear regression models to evaluate the influence of diverse structural features on merger and acquisition activities. In [26], Lee and Geum classified merger and acquisition patterns among startups using data analysis techniques like K-means clustering and visualization. They identified five distinct clusters for M&A activities within the Crunchbase dataset. The authors of [27] integrated Crunchbase data on mergers and acquisitions with other sources such as Zephyr, predicting future acquisitions through the application of machine learning and network algorithms.
@@
To the best of our knowledge, our work is the first study on the network perspective of acquisitions. Additionally, unlike prior studies, we do not confine our focus to a specific time bracket, leading to a more holistic view of the networks.
@@ 3 Methods and materials
In this section, we provide a descriptive explanation of the data and the metrics that we utilized in this study.
@@ 3.1 Data
In this part, we investigate the dataset analyzed in this paper. Figure 1 illustrates a simplified entity relationship diagram showcasing the Crunchbase data collected from the Crunchbase website by the end of July 2023. Crunchbase is a company that provides information about acquisitions, investments, and funding activities of startups. Within this comprehensive dataset, our focus was on three specific tables: organizations, acquisitions, and organization descriptions. These tables were merged based on their unique identifiers.
@@
Figure 1: Entity relationship diagram of the data.
@@ 3.1.1 Data preprocessing
In this part, we outline the steps taken to prepare the data for our analyses. First, we remove acquisitions where the date is unspecified or the type is not labeled as an acquisition. Next, we eliminate duplicate acquisitions, ensuring that only one acquisition remains for each pair of acquirer and acquiree on a specific date. Subsequently, we exclude organizations from the acquisitions if details such as their country, region, city, description, category groups list, category list, or founding date are unspecified. To ensure features are in a single-valued format, columns with list values are converted into singular values. To achieve this, we preprocess the descriptions and categories provided for each organization. We then calculate the cosine similarity between the description and each category to identify the primary category of that organization. The same process is applied to category groups. It is important to note that the categories and category groups in the initial data were sorted alphabetically, making it impossible to infer the primary ones without utilizing NLP techniques. Finally, date columns are transformed into the year-month format and then converted into numerical format for comparison across different dates. After these steps, we acquired complete data on 92,788 acquisitions, involving 119,263 organizations, out of the initial 152,923 acquisitions present in the raw data. In essence, 39.32% of the initial data was removed during the data cleansing stage.
@@ 3.1.2 Data exploration
In this part, we analyze the distribution of various features within our dataset to gain deeper insights into their characteristics.
@@
Figure 2: Representation of company distribution across countries on a world map, with the size of each country (depicted as circles) proportional to its frequency.
Figure 3 presents the distribution of companies across various cities worldwide. As per this figure, London, New York, and San Francisco host the highest number of startups, followed by Paris and Toronto in the fourth and fifth positions, respectively. These observations are consistent with prior studies highlighting London and Paris as significant European hubs [31], alongside New York and San Francisco as prominent cities for building startups [32, 33].
Figure 3: Representation of company distribution across cities on a world map, with the size of each city (depicted as circles) proportional to its frequency.
@@
Figure 4: Bubble plot of companies’ primary categories.
Figure 5 presents the distribution of companies based on their founding years, spanning from 1900. The distribution demonstrates a leftward skew, suggesting a gradual rise in startups from 1900 to 1966, succeeded by a rapid increase leading to its initial peak in 1999. This surge may be associated with the dot-com boom in the late 1990s, contributing to a rise in technology-driven startups [34]. Following this, there is a distinct decrease in 2002, succeeded by a resurgence leading to a second peak in 2012. However, post-2014, there is a discernible decline. The decrease in 2002 could be related to the aftermath of the dot-com boom, which resulted in financial problems for internet-based companies [35].
Figure 5: Distribution of companies across founding years.
Figure 6 illustrates the distribution of acquisitions across different years, beginning from 1908 (the earliest transaction date in our dataset). The distribution exhibits a left-skewed pattern, with a gradual increase in acquisitions followed by a steep upward trend starting from 2000. This increasing pattern could also be due to the dot-com crash. Ultimately, the highest peak in 2021 can be linked to businesses recovering post the COVID-19 pandemic, resulting in a rise in both acquisitions and investments [36].
Figure 6: Distribution of acquisitions across different years.
@@ 3.2 Metrics
In this part, we provide the definitions of the metrics we calculate in different parts of our analyses.
@@ 3.2.1 Transitivity
The transitivity of a network is calculated as the proportion of all potential triangles existing within the network. The formula for transitivity is:
@@
as proved by [37]. We utilize this metric to assess the strength of the community structures of the acquisition network and the common acquirer network.
@@ 3.2.2 Average clustering coefficient
The average clustering coefficient of a network indicates the tendency of nodes to cluster together. The formula for the average clustering coefficient is as follows:
@@
where $n$ is the number of nodes in the network, and $c_{v}$ is the local clustering coefficient of node $v$ [38]. We utilize this metric to assess the level of interconnectedness and community structure within the acquisition network and the common acquirer network.
@@ 3.2.3 Average shortest path length
The average shortest path length of a network is computed as the average length of the shortest paths between all pairs of nodes in the network. The formula for average shortest path length is as follows:
@@
where $d(v,u)$ is the length of shortest path between nodes $v$ and $u$, and $n$ is the number of nodes in the network [37]. We calculate this metric for the largest connected component of the acquisition network and the common acquirer network to evaluate the mean distance between each pair of organizations.
@@ 3.2.4 Assortativity
The assortativity coefficient of a network measures the proportion of connections in the network that occur between nodes of the same type based on a specific attribute [37]. We compute this metric for various node attributes to identify which attributes are more influential in shaping the connections within the acquisition network and the common acquirer network.
@@ 3.2.5 Betweenness centrality
The betweenness centrality calculates the extent to which a node lies on paths between other nodes in a network. The formula for betweenness centrality is as follows:
@@
where $\sigma(i,u,j)$ represents the number of shortest paths between nodes $i$ and $j$ that pass through node or edge $u$, and $\sigma(i,j)$ is the total number of shortest paths between $i$ and $j$. The sum is computed over all pairs $i$, $j$ of distinct nodes in the network [37]. We calculate this measure for the nodes in the cross-city acquisition network to determine which cities are influential in terms of their control over acquisition transactions between other cities.

### v3

# Understanding Trends, Patterns, and Dynamics in Global Company Acquisitions: A Network Perspective

@@ Abstract
Studying acquisitions offers invaluable insights into startup trends, aiding informed investment decisions for businesses. However, the scarcity of studies in this domain prompts our focus on shedding light in this area. Employing Crunchbase data, our study delves into the global network of company acquisitions using diverse network analysis techniques. Our findings unveil an acquisition network characterized by a primarily sparse structure comprising localized dense connections. We reveal a prevalent tendency among organizations to acquire companies within their own country and industry, as well as those within the same age bracket. Furthermore, we show that the country, region, city, and category of the companies can affect the formation of acquisition relationships between them. Our temporal analysis indicates a growth in the number of weakly connected components of the network over time, accompanied by a trend toward a sparser network. Through centrality metrics computation in the cross-city acquisition network, we identify New York, London, and San Francisco as pivotal and central hubs in the global economic landscape. Finally, we show that the United States, United Kingdom, and Germany are predominant countries in international acquisitions. The insights from our research assist policymakers in crafting better regulations to foster global economic growth, and aid businesses in deciding which startups to acquire and which markets to target for expansion.
@@ 1 Introduction
Crunchbase [^1] https://www.crunchbase.com, PitchBook [^2] https://pitchbook.com, and VentureSource [^3] https://www.venture-source.com are among the most popular platforms providing extensive datasets, collecting information about global firms and their transactions, making them valuable data sources for evaluating acquisitions [57]. However, Crunchbase stands out as the most complete and comprehensive database among them, covering the greatest number of entries and features [25]. The Crunchbase team gathers data through various channels, including monthly portfolio updates from a global network of large investment firms, as well as contributions from executives, entrepreneurs, and investors who actively update company profile pages on Crunchbase. Additionally, they utilize artificial intelligence, machine learning algorithms, and data analysts to verify the accuracy of the data both automatically and manually [^4] https://support.crunchbase.com/hc/en-us/articles/360009616013-Where-does-Crunchbase-get-their-data. Crunchbase data provides information about companies and their transactions from all around the world, covering a wide range of industries from technology and life sciences to finance and consumer goods. Furthermore, this platform offers detailed information regarding acquisitions, including type, date, and price, making it a useful data source for analyzing various facets of these transactions. Therefore, we leverage this insightful data for our research to provide a holistic view of company acquisitions.
Social network analysis (SNA) has many applications, ranging from assessing robustness in complex networks to examining signal propagation within them [2, 37]. In this study, we use SNA to construct and investigate the acquisition network of companies, where nodes represent the companies involved in acquisition transactions, and directed edges represent the connections from acquirer companies to acquirees. Then we explore the projections of this network onto acquiring and acquired firms, cities, and countries using Crunchbase data. We employ a network science approach in our analyses and interpret various structural features of these networks to gain a deeper understanding of dynamics and trends in acquisitions over time.
@@
  • RQ1: What features influence the interconnections between organizations within the acquisition network?
  • RQ2: How has the structure of the acquisition network evolved over time?
  • RQ3: Which cities or countries serve as global economic hubs in terms of acquisitions?
@@
    We calculate and interpret structural features such as transitivity, average clustering coefficient, and the size of the largest connected components for these networks.
@@
The rest of the paper is structured as follows. In the ”Related work” section, we explore previous studies based on Crunchbase data and juxtapose their scope and perspectives with ours. In the ”Methods and materials” section, we provide an explanation of our data and its features, along with the definition of the constructed network and metrics that we calculate in this study. In the ”Results” section, we describe our analyses and findings. Following this, we answer our research questions and discuss the implications, limitations, and future directions of our study in the ”Discussion” section. Finally, we conclude the paper in the ”Conclusion” section.
@@ 2 Related work
Different economic and managerial studies have utilized Crunchbase data since its foundation in 2007 to gain insights into company interactions, market trends, and strategic behaviors [18, 5]. They have employed various perspectives, focused on different sectors, geographic areas and time brackets, and utilized a wide range of data analysis techniques to address their research inquiries. Several studies have utilized Crunchbase data to forecast firms’ success using various methodologies, including machine learning and network analysis, with different sets of features such as demographic, geographical, and fundamental company information obtained from Crunchbase or LinkedIn, as well as industry-specific features as predictors [66, 63, 40, 9]. Some studies employed network analysis techniques to examine investor funding or investment behavior [47, 58, 67, 35]. Others have concentrated on analyzing market trends within specific industries like software, healthcare, manufacturing, and FinTech [16, 32, 26, 20, 14].
@@
However, only a few studies have sought to examine mergers and acquisition patterns using Crunchbase data. In [19], the authors constructed temporal networks representing international mergers and acquisitions from this dataset, analyzing both the topological and geographical characteristics of these networks. Additionally, they conducted a comparative analysis between these networks and the international trade network, aiming to discern differences and similarities, thus gaining valuable insights. Guo et al. [31] concentrated on Chinese companies while investigating merger and acquisition networks. They computed various centrality measures to scrutinize M&A behaviors and constructed multiple linear regression models to evaluate the influence of diverse structural features on merger and acquisition activities. In [44], Lee and Geum classified merger and acquisition patterns among startups using data analysis techniques like K-means clustering and visualization. They identified five distinct clusters for M&A activities within the Crunchbase dataset. The authors of [1] integrated Crunchbase data on mergers and acquisitions with other sources such as Zephyr, predicting future acquisitions through the application of machine learning and network algorithms.
@@
To the best of our knowledge, our work is the first study on the network perspective of acquisitions in which acquiring firms and startups are analyzed geographically, determining economic hubs around the world. Additionally, unlike prior studies, we do not confine our focus to a specific time bracket, leading to a more holistic view of the acquisition patterns.
@@ 3.1 Data
Figure 1 illustrates a simplified entity relationship diagram showcasing the Crunchbase data collected from the Crunchbase website by the end of July 2023. Crunchbase is a company that provides information about acquisitions, investments, and funding activities of startups. Within this comprehensive dataset, our focus was on three specific tables: organizations, acquisitions, and organization descriptions. These tables were merged based on their unique identifiers.
@@
Figure 1: Entity-relationship diagram of the data, comprising three tables: organizations, acquisitions, and organization descriptions. This figure shows the relationships between these entities and how they interact within the database.
@@ 3.1.1 Data preprocessing
To prepare the data for our analyses, we take the follwing steps. First, we remove acquisitions where the date is unspecified or the type is not labeled as an acquisition. Next, we eliminate duplicate acquisitions, ensuring that only one acquisition remains for each pair of acquirer and acquiree on a specific date. Subsequently, we exclude organizations from the acquisitions if details such as their country, region, city, description, category groups list, category list, or founding date are unspecified. To ensure features are in a single-valued format, columns with list values are converted into singular values. To achieve this, we preprocess the descriptions and categories provided for each organization. We then calculate the cosine similarity between the description and each category to identify the primary category of that organization. The same process is applied to category groups. It is important to note that the categories and category groups in the initial data were sorted alphabetically, making it impossible to infer the primary ones without utilizing NLP techniques. Finally, date columns are transformed into the year-month format and then converted into numerical format for comparison across different dates. After these steps, we acquired complete data on 92,788 acquisitions, involving 119,263 organizations, out of the initial 152,923 acquisitions present in the raw data. In essence, 39.32% of the initial data was removed during the data cleansing stage.
@@ 3.1.2 Data exploration
Figure 2: Representation of company distribution across countries on a world map. Since the frequencies of companies follow a power law distribution, we used a logarithmic transformation to ensure that significant countries are highlighted by their colors. This figure illustrates that Western countries have a much greater number of startups compared to the rest of the world.
@@
Figure 3: Bubble plot of companies’ primary categories. The size of each bubble and its label corresponds to the frequency of its category. This figure demonstrates that Software, Healthcare, and Manufacturing are the prominent categories among startups.
Figure 4 illustrates the frequency of companies’ founding years and acquisition years, spanning from 1990. The distribution of companies based on their founding years demonstrates an increasing pattern, leading to a significant initial peak in 1999. This surge may be associated with the dot-com boom in the late 1990s, contributing to a rise in technology-driven startups [6]. Following this, there is a distinct decrease in 2002, succeeded by a resurgence leading to a second peak in 2012. However, post-2014, there is a discernible decline. The decrease in 2002 could be related to the aftermath of the dot-com boom, which resulted in financial problems for internet-based companies [36].
According to Figure 4, the distribution of acquisitions across different years exhibits a left-skewed pattern, with a gradual increase in acquisitions followed by a steep upward trend starting from 2000. This increasing pattern could also be due to the dot-com crash. Ultimately, the highest peak in 2021 can be linked to businesses recovering post the COVID-19 pandemic, resulting in a rise in both acquisitions and investments [43].
Figure 4: Distributions of companies’ founding years and acquisitions across different years starting from 1990. This figure illustrates that these distributions are affected by worldwide events such as the dot-com boom and the COVID-19 pandemic.
3.2 Building networks
To illustrate how we construct each network, we consider a sample of the data provided in Figure 5.
Figure 5: An example of data comprising companies’ countries, cities, and their acquisition relationships.
3.2.1 Acquisition network
In this network, nodes represent companies engaged in acquisitions. We establish directed edges from company $u$ to company $v$ if $u$ has acquired $v$, resulting in a directed network.
3.2.2 Common acquirer network
To construct this network, we establish an edge between two companies whenever a common entity has acquired both of them. This results in an undirected weighted network, where the weight of an edge signifies the number of shared acquirers between two companies.
3.3 Common acquiree network
To build this network, we establish an undirected edge between two companies if they have acquired the same company. The weight of these edges corresponds to the number of shared acquirees between the two companies. This process results in an undirected weighted network.
Figure 6 represents examples of the acquisition network, common acquirer network, and common acquiree network according to the data in Figure 5.
Figure 6: Networks based on the data in Figure 5: a Acquisition network, b Common acquirer network, c Common acquiree network
3.3.1 Cross-city acquisition network
In the cross-city acquisition network, nodes represent cities of companies, and directed edges connect city $u$ to city $v$ if a company from city $u$ has acquired a company in city $v$. The weight of each edge represents the number of acquisitions between the two cities.
3.3.2 Cross-border acquisition network
To construct the cross-border acquisition network, we establish a directed edge from country $u$ to country $v$ whenever a company in $u$ acquires a company in $v$. This creates a weighted directed network where edge weights signify the number of acquisitions between countries.
Figure 7 shows examples of the cross-city acquisition network and cross-border acquisition network, derived from the data in Figure 5.
Figure 7: Networks based on the data in Figure 5: a Cross-city acquisition network, b Cross-border acquisition network
@@ 3.4.1 Transitivity
We measure the transitivity to assess the strength of the community structures of the acquisition network and its projections. This metric aligns with our research as it identifies the level of cohesion and resilience in the network, which are useful for making acquisition decisions. The transitivity of a network is calculated as the proportion of all potential triangles existing within the network. The formula for transitivity is:
@@
as proved by [50].
@@ 3.4.2 Average clustering coefficient
We measure the average clustering coefficient to assess the level of interconnectedness and community structure within the acquisition network and its projections. This metric is relevant to our study as it specifies the degree to which companies tend to create tightly-knit groups within the network, providing a better understanding of market dynamics. The formula for the average clustering coefficient is as follows:
@@
where $n$ is the number of nodes in the network, and $c_{v}$ is the local clustering coefficient of node $v$ [38].
@@ 3.4.3 Average shortest path length
We measure the average shortest path length for the largest connected component of the acquisition network and its projections to evaluate the mean distance between each pair of companies. This metric is useful for our research as it evaluates the connectivity of the network and the speed at which market trends and economic influences can propagate through it. The average shortest path length of a network is computed as the average length of the shortest paths between all pairs of nodes in the network. The formula for average shortest path length is as follows:
@@
where $d(v,u)$ is the length of shortest path between nodes $v$ and $u$, and $n$ is the number of nodes in the network [50].
@@ 3.4.4 Assortativity
We measure the assortativity for various node attributes to identify which attributes are more influential in shaping the connections within the acquisition network and its projections. This metric captures the tendency among companies to engage in acquisition transactions with other companies that have a similar set of attributes, making it fitting for our research. The assortativity coefficient of a network calculates the proportion of connections in the network that occur between nodes of the same type based on a specific attribute [50].
@@ 3.4.5 Betweenness centrality
We measure the betweenness centrality for the nodes in the cross-city acquisition network to determine which cities are influential in terms of their control over acquisition transactions between other cities. Betweenness centrality is helpful for this purpose as it identifies cities that act as intermediaries between different cities within the network, which can be targeted for strategic acquisitions. This metric calculates the extent to which a node lies on paths between other nodes in a network. The formula for betweenness centrality is as follows:
@@
where $\sigma(i,u,j)$ represents the number of shortest paths between nodes $i$ and $j$ that pass through node or edge $u$, and $\sigma(i,j)$ is the total number of shortest paths between $i$ and $j$. The sum is computed over all pairs $i$, $j$ of distinct nodes in the network [50].